### v1

# An effective van der Corput-type method in higher dimensions

Abstract.
[…]
1. Introduction
For the oscillatory integral
(1.1) $I(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda f(x)}\varphi(x)\,\mathrm{d}x.$
The classical van der Corput lemma states that
Lemma 1.1.
If $n=1,f^{(k)}(x)\geq 1$ on $[a,b]$, and in the case $k=1$, $f^{{}^{\prime}}(x)$ is also monotone, then there is a constant $C_{\varphi}$ such that
(1.2) $|I(\lambda)|\leq C_{\varphi}\lambda^{-1/k}$
The importance of van der Corput lemma lies in the fact that it not only provides the optimal decay rate but also yields uniform estimates. However one can not generally expect both optimal and uniform decay estimates for oscillatory integrals in higher dimensions. There is a higher-dimensional analogue of van der Corput lemma
Lemma 1.2.
If $\frac{\partial^{\alpha}f}{\partial x^{\alpha}}\geq 1$ holds on the support of $\varphi$ for some multi index $\alpha=(\alpha_{1},\cdots,\alpha_{2})$, then
(1.3) $\left|I(\lambda)\right|\leq C_{\varphi}\lambda^{-1/|\alpha|},$
where the constant also depends on higher derivatives of $f$.
It should be noted that the decay estimate above is, in general, neither optimal nor uniform. Optimal and uniform estimates are each important to analysts, albeit from different perspectives. For uniform estimates we refer the readers to [CW99] and [BGZZK24].
If one focuses only on the optimal decay of an oscillatory integral, the problem is inherently local. In the local setting, more refined information, such as asymptotic estimates or even full asymptotic expansions, may be expected. The stationary phase principle shows that the behavior of an oscillatory integral is governed by the critical points of its phase: well-behaved critical points yield better decay, while poorly behaved ones lead to weaker decay. In the nondegenerate case, the sharp decay follows from the classical stationary phase method. For phase functions of algebraic type, such as polynomials or real-analytic functions, critical points correspond to singularities of hypersurfaces, motivating the use of algebraic-geometric techniques, in particular resolution of singularities, in analysis.
In the 1960s, Arnold proposed the following hypothesis:
For oscillatory integrals with real-analytic phase functions, the principal term of the asymptotic expansion is determined by the Newton distance of the phases.
Later, Varchenko confirmed the hypothesis for generic phases in the seminal work [Var76], using toric resolution of singularities.
For the reader’s convenience, we briefly recall Varchenko’s approach. One first rewrites $I(\lambda)$ in the form
(1.4) $I(\lambda)$ $=\int_{0}^{+\infty}e^{i\lambda t}J(t)\,\mathrm{d}t+\int_{-\infty}^{0}e^{i\lambda t}J(t)\,\mathrm{d}t$
(1.5) $=\int_{0}^{+\infty}e^{i\lambda t}J(t)\,\mathrm{d}t+\int_{0}^{+\infty}e^{-i\lambda t}J(-t)\,\mathrm{d}t$
where $J(t)$ is the Gelfand-Leray function
(1.6) $J(t)=\int_{f=t}\varphi\cdot\omega,\quad t>0,$
and $\omega$ is the Gelfand-Leray form determined by
(1.7) $\omega\wedge\,\mathrm{d}t=\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n}.$
Once the asymptotic expansion of $J(t)$ is known, the corresponding expansion for $I(\lambda)$ follows immediately. The function $J(t)$ is connected to the local zeta function
(1.8) $Z_{f}(s)$ $=\int_{\mathbb{R}^{n}}\left|f(x)\right|^{s}\varphi(x)\,\mathrm{d}x$
(1.9) $=\int_{f\geq 0}(f(x))^{s}\varphi(x)\,\mathrm{d}x+\int_{f<0}(-f(x))^{s}\varphi(x)\,\mathrm{d}x$
(1.10) $:=Z_{f}^{+}(s)+Z_{f}^{-}(s)$
through inverse Mellin’s transform as follows
(1.11) $J(t)=\int_{c-i\infty}^{c+i\infty}t^{-s-1}Z_{f}^{+}(s)ds,\quad J(-t)=\int_{c-i\infty}^{c+i\infty}t^{-s-1}Z_{f}^{-}(s)ds.$
This reduces the study of the study of $J(t)$ to the meromorphic continuation of $Z_{f}(s)$ to all of $\mathbb{C}$. This continuation was conjectured by Gelfand in 1954, and later proved independently by Bernstein-Gelfand [BG69] and by Atiyah [Ati70]. Bernstein [Ber72] subsequently gave another proof using what is now known as Bernstein-Sato polynomial, introduced in the context of $\mathcal{D}$-module theory. The proofs of both Bernstein-Gelfand and Atiyah rely on Hironaka’s celebrated resolution theorem [Hir64]. However, Hironaka’s result is purely existential, whereas Arnold’s hypothesis points directly to an explicitly computable quantity-the Newton distance of the phase function. This is precisely what Varchenko accomplished: for generic phase functions, he employed a special type of resolution to relate the principal term of the asymptotic expansion to the geometry of the Newton polyhedron. More explicitly, Varchenko obtained the asymptotic expansion
(1.12) $I(\lambda)\approx\sum_{k}\sum_{l=1}^{l_{k}}a_{k,l}\lambda^{-p_{k}}\ln^{l-1}\lambda,$
where $\{p_{k}\}$ forms an arithmetic progression with the leading exponent $p_{1}$ determined by the Newton distance of the phase.
This leads to a somewhat awkward situation: in most applications, analysts are primarily interested in the principal term, yet obtaining it requires digesting Varchenko’s full proof to access the stronger asymptotics. If one wishes to avoid algebraic geometry, it becomes necessary to develop analytic tools that replicate the effect of toric resolution. In this direction, Greenblatt made substantial contributions: by introducing a remarkable resolution algorithm in [Gre08], he obtained asymptotic expansions for a broad class of smooth phase functions; see also [Gre09, Gre10a, Gre10b]. In a related development, Gilula [Gil18] established analogous results using a purely real-variable approach. Both proofs rely on a delicate decomposition of the domain, applying van der Corput lemma locally and summing the contributions of each piece. Pursuing a purely analytic proof, however, comes at the cost of losing much of the geometric intuition that algebraic methods naturally provide. Further advances on the asymptotic behavior of oscillatory integrals and on the meromorphic continuation of local zeta functions for general smooth functions were achieved in [CKN13, KN15, KN16b, KN16a, KN20].
In this work, we aim to strike an optimal balance: reducing analytic complexity while retaining the key geometric intuition.
Before presenting our main novelty and key observation, we revisit the aforementioned higher-dimensional analogue of van der Corput lemma. We briefly outline its proof, discuss its strengths and limitations, and then introduce the key improvements underlying our approach. Consider
(1.13) $I(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda f(x)}\varphi(x)\,\mathrm{d}x,$
where $\varphi$ is a smooth cut-off, and assume that
(1.14) $\frac{\partial^{\alpha}f}{\partial x^{\alpha}}\geq 1.$
Following Stein[SM93], there exists a finite collection of unit vectors $\{\xi_{j}\}$ such that
(1.15) $\partial^{\alpha}_{x}f(x)=\sum_{j}a_{j}\left(\xi_{j}\cdot\nabla\right)^{\left|\alpha\right|}f(x).$
Hence the lower bound on $\partial^{\alpha}_{x}f(x)$ implies that for some $\xi_{j}$ we have $\left(\xi_{j}\cdot\nabla\right)^{\left|\alpha\right|}f(x)>c$ for an absolute constant $c>0$. By rotating coordinates so that this $\xi_{j}$ becomes a coordinate axis, one may apply the one-dimensional van der Corput lemma in that direction and obtain the desired decay.
This argument has two clear advantages. First, the rotation is an invertible linear transformation, so the Jacobian plays no role. Second, the directional derivatives are collected into a single direction, allowing direct use of the one-dimensional lemma. Its limitation, however, stems from the same linearity: although we obtain a direction along which to apply van der Corput lemma, we have no precise control over this direction, and in general it does not yield the optimal decay. This reflects the broader fact that linear changes of variables do not reveal sufficient information about the singularities of the phase. By contrast, nonlinear transformations, such as blow-ups arising in algebraic geometry, capture precisely the singular structure that must be resolved. So our key observation is
Applying a toric resolution of singularities to the phase function reveals explicit geometric directions adapted to its singularities, and these directions are exactly those along which one-dimensional van der Corput estimates produce the optimal decay rate of the oscillatory integral.
2. A Motivating Example and Generalizations
In this section we begin with a toy model that illustrates the central idea, and then extend the argument to the general case.
Example 2.1.
Consider the toy model $f(x_{1},x_{2})=x_{1}^{2}x_{2}$, applying van der Corput lemma along the direction of $\xi_{j}$ above yields
(2.1) $|I(\lambda)|\leq C_{\varphi}\lambda^{-\frac{1}{3}}.$
However, the phase is more degenerate in the $x_{1}$-direction. If we first integrate in $x_{1}$ and apply the van der Corput lemma to $x_{1}\to x_{1}^{2}x_{2}$, we obtain
(2.2) $\left|\int_{\mathbb{R}}e^{i\lambda x_{1}^{2}x_{2}}\varphi(x_{1},x_{2})\,\mathrm{d}x_{1}\right|\leq C_{\varphi}\left|\lambda x_{2}\right|^{-\frac{1}{2}},$
and therefore
(2.3) $\left|I(\lambda)\right|\leq C_{\varphi}\lambda^{-\frac{1}{2}}.$
Thus, choosing the direction that reflects the actual singular structure of the phase yields a strictly better decay than that obtained from the linear method. Now consider the monomial
(2.4) $f(x)=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}},\quad\alpha\in\mathbb{Z}_{+}^{n}$
For such a phase, it is clear that the most singular behavior occurs along those coordinate directions $x_{j}$ for which $\alpha_{j}$ is largest; these are precisely the directions in which the phase degenerates most severely. We now state a lemma that provides the decay rate for oscillatory integrals with this monomial phase.
Lemma 2.2.
Consider
(2.5) $I(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n},$
where $\alpha,\beta\in\mathbb{Z}_{+}^{n}$, $\operatorname*{supp}\varphi\subset[-1,1]^{n}$ and denote by
(2.6) $d=\max_{j}\alpha_{j},\quad M=\#\{j:\alpha_{j}=d\},$
then it follows
(2.7) $\left|I(\lambda)\right|\leq C_{\varphi}\lambda^{-\frac{1}{d}}\ln^{M-1}\lambda.$
Proof.
Without loss of generality, we may restrict the domain of integration to the first orthant $\mathbb{R}_{+}^{n}$ and set
(2.8) $I^{+}(\lambda)=\int_{\mathbb{R}_{+}^{n}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n}.$
To obtain an estimate for $I(\lambda)$, it suffices to prove a similar estimate for $I^{+}(\lambda)$. To analyze $I^{+}(\lambda)$, we decompose the domain depending on whether the oscillation is effective or negligible. This yields a splitting
(2.9) $I^{+}(\lambda)$ $=\int_{\left\{x\in\mathbb{R}_{+}^{n}:x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}<1/\lambda\right\}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n}$
(2.10) $+\int_{\left\{x\in\mathbb{R}_{+}^{n}:x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}\geq 1/\lambda\right\}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n}$
(2.11) $:=I_{1}^{+}(\lambda)+I_{2}^{+}(\lambda),$
where $I_{1}^{+}(\lambda)$ is the contribution from a sublevel region (handled by size estimates), and $I_{2}^{+}(\lambda)$ is the contribution from the oscillatory region (handled by van der Corput-type estimates). Therefore
(2.12) $\left|I_{1}^{+}(\lambda)\right|$ $\leq\|\varphi\|_{L^{\infty}}\left|\left\{x\in[0,1]^{n}:x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}<1/\lambda\right\}\right|$
(2.13) $\leq C_{\varphi}\lambda^{-\frac{1}{d}}\ln^{M-1}\lambda.$
The last inequality follows from Lemma 3.1 of [Gre10a]. We now proceed to treat the term $I_{2}^{+}$. Without loss of generality, we may assume $\alpha_{1}=d$. Observe that
(2.14) $\left\{x\in[0,1]^{n}:x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}\geq 1/\lambda\right\}\subset\left\{x\in[0,1]^{n}:x_{j}\geq\lambda^{-\frac{1}{\alpha_{j}}},j=1,\cdots,n\right\}$
and that
(2.15) $\partial_{x_{1}}^{d}(x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}})=d!x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}.$
Applying van der Corput lemma in the $x_{1}$-variable yields
(2.16) $\left|I_{2}^{+}\right|$ $\leq C_{\varphi}\int_{\lambda^{-\frac{1}{\alpha_{n}}}}^{1}\cdots\int_{\lambda^{-\frac{2}{\alpha_{2}}}}^{1}\left(d!\lambda x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}\right)^{-\frac{1}{d}}\,\mathrm{d}x_{2}\cdots\,\mathrm{d}x_{n}$
(2.17) $\leq C_{\varphi}\lambda^{-\frac{1}{d}}\ln^{M-1}\lambda.$
Now we state a more general lemma.
Lemma 2.3.
Consider the oscillatory integral
(2.18) $I(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\left|x_{1}^{\beta_{1}}\cdots x_{n}^{\beta_{n}}\right|\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n},$
where $\alpha,\beta\in\mathbb{Z}_{+}^{n}$ and $\operatorname*{supp}\varphi\subset[-1,1]^{n}$. Set
(2.19) $d=\max_{j}\frac{\alpha_{j}}{\beta_{j}+1},\quad M=\#\left\{j:\frac{\alpha_{j}}{\beta_{j}+1}=d\right\}.$
Then one has the decay estimate
(2.20) $\left|I(\lambda)\right|\leq C_{\varphi}\lambda^{-\frac{1}{d}}\ln^{M-1}\lambda.$
Proof.
We first treat the case $d\geq 1$. The case $d<1$ can be reduced to this one by integration-by-parts argument. Similar to the preceding lemma, the estimate of $I(\lambda)$ reduces to estimating $I^{+}(\lambda)$ and $I^{-}(\lambda)$. We write
(2.21) $I_{1}^{+}(\lambda)$ $=\int_{\left\{x\in\mathbb{R}_{+}^{n}:x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}<1/\lambda\right\}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\cdots x_{n}^{\beta_{n}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n},$
(2.22) $I_{2}^{+}(\lambda)$ $=\int_{\left\{x\in\mathbb{R}_{+}^{n}:x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}\geq 1/\lambda\right\}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\cdots x_{n}^{\beta_{n}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n}.$
To estimate $I_{1}^{+}(\lambda)$, we perform the change of variables
(2.24) $x_{j}=y_{j}^{\frac{1}{\beta_{j}+1}},\quad j=1,\cdots,n.$
Then
(2.25) $\left|I_{1}^{+}(\lambda)\right|\leq C_{\varphi}\int_{\left\{x\in[0,1]^{n}:y_{1}^{\frac{\alpha_{1}}{\beta_{1}+1}}\cdots y_{n}^{\frac{\alpha_{n}}{\beta_{n}+1}}<1/\lambda\right\}}\,\mathrm{d}y_{1}\cdots\,\mathrm{d}y_{n},$
and by Lemma 3.1 in [Gre10a], we obtain
(2.26) $\left|I_{1}^{+}(\lambda)\right|\leq C_{\varphi}\lambda^{-\frac{1}{d}}\ln^{M-1}\lambda.$
We now deal with $I_{2}^{+}(\lambda)$. Without loss of generality we assume that $\frac{\alpha_{1}}{\beta_{1}+1}=d$. Observe that
(2.27) $I_{2}^{+}(\lambda)$ $=\int_{\left\{x\in\mathbb{R}_{+}^{n}:x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}\geq 1/\lambda\right\}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\cdots x_{n}^{\beta_{n}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n}$
(2.28) $=\int_{D(x_{2},\cdots,x_{n})}\left[\int_{L(x_{2},\cdots,x_{n})}^{1}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\cdots x_{n}^{\beta_{n}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\right]\,\mathrm{d}x_{2}\cdots\,\mathrm{d}x_{n}$
(2.29) $\leq\int_{\lambda^{-\frac{1}{\alpha_{n}}}}^{1}\cdots\int_{\lambda^{-\frac{2}{\alpha_{2}}}}^{1}\left|\int_{L(x_{2},\cdots,x_{n})}^{1}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\cdots x_{n}^{\beta_{n}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\right|\,\mathrm{d}x_{2}\cdots\,\mathrm{d}x_{n},$
where $L(x_{2},\cdots,x_{n})=\left(\lambda x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}\right)^{-\frac{1}{\alpha_{1}}}$ and the outer domain $D(x_{2},\cdots,x_{n})$ is contained in the box $\lambda^{-\frac{1}{\alpha_{2}}},1]\times\cdots\times[\lambda^{-\frac{1}{\alpha_{n}}},1]$.
By integration by parts, we can see that to estimate the inner integral, it suffices to estimate
(2.30) $\int_{L(x_{2},\cdots,x_{n})}^{1}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\,\mathrm{d}x_{1}.$
A direct computation yields
(2.31) $\int_{L(x_{2},\cdots,x_{n})}^{1}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\,\mathrm{d}x_{1}$
(2.32) $=\left.\frac{e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}}{i\alpha_{1}\lambda x_{1}^{\alpha_{1}-1}\cdots x_{n}^{\alpha_{n}}}\right|_{L(x_{2},\cdots,x_{n})}^{1}-\frac{\beta_{1}+1-\alpha_{1}}{i\alpha_{1}\lambda x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}}\int_{L(x_{2},\cdots,x_{n})}^{1}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}-\alpha_{1}}\,\mathrm{d}x_{1}.$
Using the van der Corput lemma for the last integral (we note that when $d=1$, the coefficient $\beta_{1}+1-\alpha_{1}=0$, and the last integral does not appear), we obtain
(2.33) $\left|\int_{L(x_{2},\cdots,x_{n})}^{1}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\,\mathrm{d}x_{1}\right|\leq C\lambda^{-1}L(x_{2},\cdots,x_{n})^{\beta_{1}+1-\alpha_{1}}\left(x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}\right)^{-1}.$
Hence
(2.34) $\left|\int_{L(x_{2},\cdots,x_{n})}^{1}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}x_{2}^{\beta_{2}}\cdots x_{n}^{\beta_{n}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\right|$
(2.35) $\leq C_{\varphi}\lambda^{-1}L(x_{2},\cdots,x_{n})^{\beta_{1}+1-\alpha_{1}}x_{2}^{\beta_{2}-\alpha_{2}}\cdots x_{n}^{\beta_{n}-\alpha_{n}}$
(2.36) $=C_{\varphi}\lambda^{-1}\left(\lambda x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}\right)^{-\frac{\beta_{1}+1-\alpha_{1}}{\alpha_{1}}}x_{2}^{\beta_{2}-\alpha_{2}}\cdots x_{n}^{\beta_{n}-\alpha_{n}}$
(2.37) $=C_{\varphi}\lambda^{-1/d}x_{2}^{\beta_{2}-\alpha_{2}/d}\cdots x_{n}^{\beta_{n}-\alpha_{n}/d}.$
Consequently,
(2.38) $I_{2}^{+}(\lambda)$ $\leq C_{\varphi}\lambda^{-1/d}\int_{\lambda^{-\frac{1}{\alpha_{n}}}}^{1}\cdots\int_{\lambda^{-\frac{2}{\alpha_{2}}}}^{1}x_{2}^{\beta_{2}-\alpha_{2}/d}\cdots x_{n}^{\beta_{n}-\alpha_{n}/d}\,\mathrm{d}x_{2}\cdots\,\mathrm{d}x_{n}$
(2.39) $\leq C_{\varphi}\lambda^{-\frac{1}{d}}.$
This completes the proof for the case $d\geq 1$.
We now consider the case $d<1$. The key point is that the higher order of vanishing of the weight $x^{\beta}$ near the origin allows us to perform repeated integration by parts without incurring any loss. Without loss of generality we assume that
(2.40) $\frac{\alpha_{1}}{\beta_{1}+1}=d.$
Choosing an integer $K>0$ such that
(2.41) $\beta_{1}+1-K\alpha_{1}\geq 0\quad\text{and}\quad\beta_{1}+1-(K+1)\alpha_{1}\leq 0.$
(2.42) $I^{+}(\lambda)$ $=\int_{\mathbb{R}_{+}^{n}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}x_{1}^{\beta_{1}}\cdots x_{n}^{\beta_{n}}\varphi(x_{1},\cdots,x_{n})\,\mathrm{d}x_{1}\cdots\,\mathrm{d}x_{n}$
(2.43) $=\int_{\mathbb{R}_{+}^{n-1}}\left[\int_{\mathbb{R}_{+}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\varphi(x_{1},\cdots,x_{n})x_{1}^{\beta_{1}}\,\mathrm{d}x_{1}\right]x_{2}^{\beta_{2}}\cdots x_{n}^{\beta_{n}}\,\mathrm{d}x_{2}\cdots\,\mathrm{d}x_{n}.$
Applying $K$ integrations by parts the resulting integrand still has nonnegative order at the boundary, while one additional integration by parts would produce a negative power of $x_{1}$. More precisely, we obtain
(2.44) $\frac{C}{\left(i\lambda x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}\right)^{K}}\int_{\mathbb{R}_{+}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\varphi(x_{1},\cdots,x_{n})x_{1}^{\beta_{1}-K\alpha_{1}}\,\mathrm{d}x_{1}.$
We also note that if $1/d$ is an integer, this final integration by parts is unnecessary; in that case, a simple change of variables suffices to obtain the desired estimate. Consequently, the integral reduces to
(2.45) $\frac{C}{\lambda^{K}}\int_{\mathbb{R}_{+}^{n}}\int_{\mathbb{R}_{+}}e^{i\lambda x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}}\varphi(x_{1},\cdots,x_{n})x_{1}^{\beta_{1}-K\alpha_{1}}\cdots x_{n}^{\beta_{n}-K\alpha_{n}}\,\mathrm{d}x_{1}\,\mathrm{d}x_{2}\cdots\,\mathrm{d}x_{n}.$
By construction, we then have
(2.46) $\frac{\alpha_{1}}{\beta_{1}+1-K\alpha_{1}}=\max_{j}\frac{\alpha_{j}}{\beta_{j}+1-K\alpha_{j}}\geq 1.$
Thus, applying the same argument as in the case $d\geq 1$, we obtain the desired upper bound.
3. Notations and the Main Theorem
In this section we introduce several notions about real-analytic functions that will be used throughout the paper and state the main result to be proved.
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is said to be real-analytic near the origin if, in some neighborhood of 0, it admits a convergent Taylor expansion
(3.1) $f(x)=\sum_{\alpha\in\mathbb{N}^{n}}a_{\alpha}x^{\alpha},\quad\alpha=(\alpha_{1},\cdots,\alpha_{n}).$
The Newton polyhedron of $f$ is defined as
(3.2) $\mathcal{N}_{f}:=\text{convex hull of}\left\{\alpha+\mathbb{R}_{+}^{n}:a_{\alpha}\neq 0\right\}.$
The associated Newton distance is defined by
(3.3) $d_{f}:=inf\left\{t:(t,\cdots,t)\in\mathcal{N}(f)\right\}.$
The compact face of $\mathcal{N}_{f}$ that contains the point $(d_{f},\cdots,d_{f})$ and has minimal possible dimension is called the principal face of $f$.
For any compact face $\gamma\subset\mathcal{N}_{f}$, the corresponding $\gamma$-part of $f$ is
(3.4) $f_{\gamma}=\sum_{\alpha\in\gamma}a_{\alpha}x^{\alpha}.$
A compact face $\gamma$ is said to be $\mathbb{R}$-nondegenerate if
(3.5) $\nabla f_{\gamma}\neq 0\quad\text{for all }x\in(\mathbb{R}\backslash 0)^{n},$
that is, the gradient of $f_{\gamma}$ does not vanish outside the coordinate hyperplanes. We say $f$ is $\mathbb{R}$-nondegenerate if all compact faces of its Newton polyhedron are $\mathbb{R}$-nondegenerate.
Now we are ready to state the main theorem.
Theorem 3.1.
Let
(3.6) $I(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda f(x)}\varphi(x)\,\mathrm{d}x,$
where $f$ is real-analytic and $\mathbb{R}$-nondegenerate near the origin, satisfying $f(0)=0,\nabla f(0)=0$. Assume that $\varphi$ is supported in a sufficiently small neighborhood of 0. If the codimension of principal face is $k$, then
(3.7) $\left|I(\lambda)\right|\leq C_{\varphi}\lambda^{-\frac{1}{d_{f}}}\ln^{k-1}\lambda,$
whenever $1/d_{f}$ is not an integer, and otherwise
(3.8) $\left|I(\lambda)\right|\leq C_{\varphi}\lambda^{-\frac{1}{d_{f}}}\ln^{k}\lambda.$
Remark 3.2.
If $d_{f}=1$, then under certain conditions, Varchenko [Var76] showed that the exponent of the logarithmic factor in the leading term is at most k1k-1, in agreement with Gilula’s upper bound [Gil18]. In contrast, our theorem above yields a slightly weaker estimate in this case, allowing for a logarithmic exponent up to $k$.
If we carry out the argument developed in this paper, we can also recover the main theorem of Gilula (2018) up to a logarithmic factor.
Theorem 3.3 (Gilula 2018 [Gil18]).
Consider the damped oscillatory integral
(3.9) $I(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda f(x)}x^{\beta}\varphi(x)\,\mathrm{d}x.$
Then
(3.10) $\left|I(\lambda)\right|\leq C_{\varphi}\lambda^{-\lfloor\beta+\bf{1}\rfloor}\ln^{d_{\beta}-1}\lambda,$
where $\lfloor\beta+\bf{1}\rfloor$ is the maximal number $c$ such that $\frac{\beta+\bf{1}}{c}$ is contained in $\mathcal{N}_{f}$ and $d_{\beta}$ is the greatest codimension over all faces of $\mathcal{N}_{f}$ containing $\frac{\beta+\bf{1}}{\lfloor\beta+\bf{1}\rfloor}$.
4. Key Facts About Toric Varieties
Although this material is standard for algebraic geometers, we include it here for the convenience of readers, especially analysts who may be less familiar with the algebraic-geometric background.
A toric variety is an object in algebraic geometry that plays a central role in making explicit the correspondence between combinatorial geometry and algebraic geometry. It provides an ideal testing ground for general theories, since many geometric and analytic phenomena can be translated into concrete combinatorial data.
Our ultimate goal is to construct a toric variety associated with the Newton polyhedron of the phase function. Before doing so, we introduce the basic objects underlying this construction: cones and fans.
In this paper, we only consider rational cones $\sigma\subset\mathbb{R}^{n}$, which are generated by finitely many elements of $\mathbb{Z}^{n}$. That is, given a collection of vectors $\{v_{1},\cdots,v_{n}\}$ in $\mathbb{Z}^{n}$, the cone they generate is
(4.1) $\sigma=\left\{\mu_{1}v_{1}+\cdots+\mu_{n}v_{n},\mu_{1},\cdots,\mu_{n}\geq 0\right\}.$
The skeletons of a cone $\sigma$ is the set of primitive vectors that generate $\sigma$; that is, the minimal nonzero integer vectors along the edges of $\sigma$.
A rational cone is called simplicial if the vectors making up its skeleton are linearly independent.
The fan $\Sigma$ in $\mathbb{R}^{n}$ is a finite collection of rational polyhedral cones satisfying the following conditions:
  • Every face of a cone in $\Sigma$ is also a cone in $\Sigma$;
  • The intersection of any two cones in $\Sigma$ is also a face of each.
For a fan $\Sigma$, the union $\left|\Sigma\right|:=\bigcup_{\sigma\in\Sigma}\sigma$ is called the support of $\Sigma$.
A fan $\Sigma$ is said to be simple if each cone in $\Sigma$ is simplicial and the skeletons of any cone can be extended to a basis of the integer lattice of the whole space.
An important fact about fan is the existence of simplicial subdivisions. Given a fan $\Sigma_{0}$ , the exists a fan $\Sigma$, called a simplicial subdivision of $\Sigma_{0}$, satisfying
  • $\Sigma_{0}$ and $\Sigma$ have the same support;
  • Each cone of $\Sigma$ is contained in some cone of $\Sigma_{0}$;
  • $\Sigma$ is simple.
Such subdivisions are fundamental in toric geometry, particularly for resolving singularities while maintaining control over the combinatorial structure.
We now define a rational map that is central to the construction of a toric variety. Let $\Sigma$ be a fan in $\mathbb{R}^{n}$, not necessarily simple. Consider an $n$-dimensional cone $\sigma\in\Sigma$ generated by skeletons consisting of vectors
(4.2) $\alpha^{1}=(\alpha_{1}^{1},\cdots,\alpha_{n}^{1}),\cdots,\alpha^{n}=(\alpha_{1}^{n},\cdots,\alpha_{n}^{n}).$
Given such a cone $\sigma$, we associate a copy of $\mathbb{R}^{n}$ denoted by $\mathbb{R}^{n}(\sigma)$. Define the map
(4.3) $\pi(\sigma):\mathbb{R}^{n}(\sigma)\to\mathbb{R}^{n}:\quad\pi(\sigma)(y_{1},\cdots,y_{n})=(x_{1},\cdots,x_{n}),$
where
(4.4) $x_{j}=y_{1}^{\alpha_{j}^{1}}\cdots y_{n}^{\alpha_{j}^{n}},\quad j=1,\cdots,n.$
The toric variety $X_{\Sigma}$ is then obtained by gluing the copies $\mathbb{R}^{n}(\sigma)$ along the images of the maps $\pi(\sigma)$ for all $\sigma\in\Sigma$. More explicitly, let $\sigma,\sigma^{\prime}\in\Sigma$ be two $n$-dimensional cones, and let $x\in\mathbb{R}^{n}(\sigma),x^{\prime}\in\mathbb{R}^{n}(\sigma^{\prime})$, we identify $x$ and $x^{\prime}$ if
(4.5) $\pi^{-1}(\sigma^{\prime})\circ\pi(\sigma):x\mapsto x^{\prime},$
that is, a point in the chart $\mathbb{R}^{n}(\sigma)$ is equivalent to a point in $\mathbb{R}^{n}(\sigma^{\prime})$ whenever the monomial map relating the two charts is defined at $x$ and maps it into $x^{\prime}$. This identification glues the charts together to form the toric variety $X_{\Sigma}$.
The key properties of toric varieties for our purposes are summarized as follows:
Facts:
  • If the fan $\Sigma$ is simple, then the associated toric variety $X_{\Sigma}$ is smooth.
  • The map $\pi:X_{\Sigma}\to\mathbb{R}^{n}$ defined on each $\mathbb{R}^{n}(\sigma)$ by $\pi(\sigma)$ is proper, that is, the preimage of a compact set is compact.
Based on the preceding facts, we are now ready to construct the smooth toric variety associated with the Newton polyhedron. Consider the real-analytic function introduced in Section 3. For any top-dimensional face $\gamma$ of its Newton polyhedron, let $\xi\in\mathbb{Z}_{+}^{n}$ denote the primitive normal vector to $\gamma$. Then $\gamma$ is characterized by the supporting hyperplane equation
(4.6) $\xi,x>=\min_{v\in V}<\xi,v>,$
where $V$ denotes the set of all vertices of $N_{f}$. We write
(4.7) $l(\xi)=\min_{v\in V}<\xi,v>$
for this minimal value. The collection of all such primitive normal vectors forms the set of generators (the skeleton) of a fan, denoted by $\Sigma_{0}$, whose support is the positive orthant $\mathbb{R}_{+}^{n}$. From the facts in the preceding section, the fan $\Sigma_{0}$ admits a simplicial subdivision $\Sigma$. Consequently, the fan $\Sigma$ determines a smooth toric variety $X_{\Sigma}$.
5. Proof of the Main Theorem
Proof.
We transform the oscillatory integral over $\mathbb{R}^{n}$ into an integral over the smooth toric variety $X_{\Sigma}$:
(5.1) $I(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda f(x)}\varphi(x)\,\mathrm{d}x=\int_{X_{\Sigma}}e^{i\lambda f\circ\pi(y)}\varphi\circ\pi(y)\left|J_{\pi}(y)\right|\,\mathrm{d}y.$
Since $\pi$ is proper, and $\varphi$ has compact support. It suffices to integrate over the restricted domain $X_{\Sigma}\bigcap\pi^{-1}(\operatorname*{supp}\varphi)$. Using a covering altlas $\{U_{k}\}$ of $X_{\Sigma}$ and a partition of unity, we can decompose $I(\lambda)$ as
(5.2) $I(\lambda)=\sum_{k}\int_{X_{\Sigma}\bigcap\pi^{-1}(\operatorname*{supp}\varphi)\bigcap U_{k}}e^{i\lambda f\circ\pi(y)}\varphi\circ\pi(y)\varphi_{k}(y)\left|J_{\pi}(y)\right|\,\mathrm{d}y.$
Let $k$ be fixed, and set
(5.3) $V=X_{\Sigma}\bigcap\pi^{-1}(\operatorname*{supp}\varphi)\bigcap U_{k},$
and for convenience, we continue to denote by $I(\lambda)$ the integral over $V$. On such a chart $U_{k}$, the coordinates are given by the rational map associated with an $n$-dimensional cone $\sigma\in\Sigma$, whose skeletons are
(5.4) $\alpha^{1}=(\alpha_{1}^{1},\cdots,\alpha_{n}^{1}),\cdots,\alpha^{n}=(\alpha_{1}^{n},\cdots,\alpha_{n}^{n}).$
A direction computation shows that the Jacobian of $\pi(\sigma)$ is
(5.5) $J_{\pi(\sigma)}(y)=\left|y_{1}\right|^{\left|\alpha^{1}\right|-1}\cdots\left|y_{n}\right|^{\left|\alpha^{n}\right|-1},\quad\left|\alpha^{j}\right|:=\alpha_{1}^{j}+\cdots+\alpha_{n}^{j}.$
We now examine the form of the phase function on the chart $U_{k}$. By construction of the skeleton vectors $\alpha^{1},\cdots,\alpha^{n}$, associated with the cone $\sigma$, the supporting hyperplane orthogonal to these vectors meets the Newton polyhedron $\mathcal{N}_{f}$ in a single vertex. Consequently, the pullback of $f$ under the associated monomial map $\pi(\sigma)$ may be written as
(5.6) $f\circ\pi(\sigma)(y)=y_{1}^{l(\alpha^{1})}\cdots y_{n}^{l(\alpha^{n})}f_{\sigma}(y_{1},\cdots,y_{n}),$
and $f(0)\neq 0$.
The exceptional set $\pi_{\sigma}^{-1}(0)$ is a union of coordinate hyperplanes:
(5.7) $\pi_{\sigma}^{-1}(0)=\bigcup_{j\in I}\{y_{j}=0\}$
where $I\subset\{1,\cdots,n\}$ collects the indices corresponding to the compact face $\gamma$ associated with $\left\{\alpha^{1},\cdots,\alpha^{s}\right\}$. Without loss of generality, we may relabel coordinates so that $I=\{1,\cdots,s\}$, and denote
(5.8) $T_{I}=\left\{(y_{1},\cdots,y_{n}):y_{j}=0,j\in I\right\}.$
Consequently, we may decompose the pullback of $f$ as
(5.9) $f\circ\pi(\sigma)(y)=y_{1}^{l(\alpha^{1})}\cdots y_{s}^{l(\alpha^{s})}y_{s+1}^{l(\alpha^{s+1})}\cdots y_{n}^{l(\alpha^{n})}\left(\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y)\right),$
where $\bar{f_{\sigma}}(0)\neq 0,p=(p_{1},\cdots,p_{n})$ satisfies $p_{j}>0$ for all $j$, and $y^{p}h(y)$ consists of terms of strictly higher order relative to the face $\gamma$. A key property of $\bar{f_{\sigma}}$ is the following. Let
(5.10) $w=(0,\cdots,0,w_{s+1},\cdots,w_{n})\in\mathbb{R}^{n}\backslash\pi(\sigma)^{-1}(0),$
if $\bar{f_{\sigma}}(w)=0$, then there exists $w_{j}\neq 0,j>s$ such that $\partial_{y_{j}}\bar{f_{\sigma}}(w)\neq 0$. This can be found in [Var76].
We proceed to estimate
(5.11) $I(\lambda)=\int_{V}e^{i\lambda y_{1}^{l(\alpha^{1})}\cdots y_{s}^{l(\alpha^{s})}y_{s+1}^{l(\alpha^{s+1})}\cdots y_{n}^{l(\alpha^{n})}\left(\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y)\right)}\left|y_{1}\right|^{\left|\alpha^{1}\right|-1}\cdots\left|y_{n}\right|^{\left|\alpha^{n}\right|-1}\varphi_{k}(y)\varphi\circ\pi(\sigma)(y)\,\mathrm{d}y$
By a partition of unity, we decompose $V$ into finitely many pieces.
(5.12) $1=\sum_{j}\psi_{j}+\sum_{l}\eta_{l}.$
Each $\psi_{j}$ is supported where $\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})$ does not change sign, while the union of the supports of the $\eta_{l}$ contains all zeroes of $\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})$. Inserting this decomposition into $I(\lambda)$ we obtain corresponding integrals $I_{j}(\lambda)$ and $I_{l}(\lambda)$. It suffices to estimate each piece separately.
On the support of $\psi_{k}$, the function $\bar{f_{\sigma}}(y)$ is bounded from zero. Thus $\bar{f_{\sigma}}$ and the higher order term may be absorbed into any one of the variables of $y_{s+1},\cdots,y_{n}$. For definiteness, we introduce
(5.13) $u_{1}=y_{1},\cdots,u_{s+1}=\epsilon y_{s+1}\left(\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y)\right),\cdots,u_{n}=y_{n}$
where $\epsilon=\pm 1$, the choice is determined by the sign of $y_{s+1}^{l(\alpha^{s+1})}$. This is an invertible change of variables provided the support of $\psi_{j}$ is chosen sufficiently small. After this substitution, the phase becomes a pure monomial, then
(5.14) $I_{j}(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda u_{1}^{l(\alpha^{1})}\cdots u_{n}^{l(\alpha^{n})}}\left|u_{1}\right|^{\left|\alpha^{1}\right|-1}\cdots\left|u_{n}\right|^{\left|\alpha^{n}\right|-1}\Phi_{j}(u)\,\mathrm{d}u.$
Applying Lemma 2.3, and notice that
(5.15) $\frac{l(\alpha^{j})}{\left|\alpha^{j}\right|}\leq d_{f},\quad j=1,\cdots,n,$
it implies
(5.16) $\left|I_{j}(\lambda)\right|\leq C_{\Phi_{j}}\lambda^{-\frac{1}{d_{f}}}\ln^{k-1}\lambda.$
Hence the required bound follows for each $I_{j}(\lambda)$
On the support of $\eta_{l}$, all zeroes lie inside, but by assumption the gradient of $\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})$ does not vanish in the support. Shrinking the support of $\eta_{l}$ if necessary, we may assume $\partial_{y_{s+1}}\bar{f_{\sigma}}\neq 0$ and $y_{s+1}\neq 0$ throughout this neighborhood. Therefore we can introduce the change of variables
(5.17) $u_{1}=y_{1},\cdots,u_{s+1}=\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y),\cdots,u_{n}=x_{n}.$
This transformation is smooth and invertible. In the new coordinates, the phase again reduces to a monomial. Specifically,
(5.18) $I_{l}(\lambda)=\int_{\mathbb{R}^{n}}e^{i\lambda u_{1}^{l(\alpha^{1})}\cdots u_{s+1}\cdots u_{n}^{l(\alpha^{n})}}\left|u_{1}\right|^{\left|\alpha^{1}\right|-1}\cdots\left|u_{s-1}^{l(\alpha^{s-1})}\right|\left|u_{s+1}^{l(\alpha^{s+1})}\right|\left|u_{n}\right|^{\left|\alpha^{n}\right|-1}\Psi_{l}(u)\,\mathrm{d}u.$
In the case $d_{f}>1$, we may directly apply Lemma 2.3 to obtain the desired upper bound. Suppose now $d_{f}\leq 1$. For this integral $I_{l}(\lambda)$, we may assume without loss of generality that
(5.19) $\frac{l(\alpha^{1})}{\left|\alpha^{1}\right|}=\max_{j}\frac{l(\alpha^{j})}{\left|\alpha^{j}\right|}.$
Fix $x_{1},\cdots,x_{s},x_{s+2},\cdots,x_{n}$, and perform $N$ integrations by parts with respect to $x_{s+1}$, choosing $N$ such that
(5.20) $\left|\alpha^{1}\right|-N\alpha_{1}\geq 0\quad\text{and}\quad\left|\alpha^{1}\right|-(N+1)\alpha_{1}\leq 0.$
then the main term of $I_{l}(\lambda)$ is
(5.21) $\frac{C}{\lambda^{N}}\int_{\mathbb{R}^{n}}e^{i\lambda u_{1}^{l(\alpha^{1})}\cdots u_{s+1}\cdots u_{n}^{l(\alpha^{n})}}\left|u_{1}\right|^{\left|\alpha^{1}\right|-1}u_{1}^{-Nl(\alpha^{1})}\cdots\left|u_{s-1}^{l(\alpha^{s-1})}\right|u_{s-1}^{-Nl(\alpha^{s-1})}\times$
(5.22) $\left|u_{s+1}^{l(\alpha^{s+1})}\right|u_{s+1}^{-Nl(\alpha^{s+1})}\cdots\left|u_{n}\right|^{\left|\alpha^{n}\right|-1}u_{n}^{-Nl(\alpha^{n})}\Psi_{l}(u)\,\mathrm{d}u.$
Observe that
(5.23) $\frac{l(\alpha^{1})}{\left|\alpha^{1}\right|-Nl(\alpha^{1})}=\max_{j}\frac{l(\alpha^{j})}{\left|\alpha^{j}\right|-Nl(\alpha^{j})}.$
Hence, an application of Lemma 2.3 once more yields
(5.24) $\left|I_{l}(\lambda)\right|\leq C_{\Phi_{j}}\lambda^{-\frac{1}{d_{f}}}\ln^{k}\lambda,$
where the additional power of the logarithm arises from the contribution of the variable $u_{s+1}$.
Summing over all pieces then completes the proof.
Acknowledgments
The author is supported by the National Natural Science Foundation of China (Grant No. 12501124) and the Jiangsu Natural Science Foundation (Grant No. BK20200308). The author thanks Prof. Xiaochun Li for sharing the insight that resolution of singularities plays a fundamental role in analysis. The author is grateful to Prof. Changxing Miao for inviting him to visit in 2023 and for encouraging him to pursue this meaningful topic. The author also thanks Prof. Toshihiro Nose for generously sharing his PhD thesis, from which the author learned many details about the rigorous analysis of toric resolution of singularities. This work was completed during a visit to Prof. Shaoming Guo at the Chern Institute of Mathematics; the author thanks Prof. Guo and the institute for their hospitality.